# 0.5 eV QCD Axion Cosmology

### v48

@@ 1 Introduction
In light of the apparent success of 0.5 eV QCD axion dark matter in resolving the Hubble tension, these astrophysical “constraints” and “exclusions” raise important questions about the conventional assumptions for a wide range of astrophysical phenomena. From the velocity dispersion of dark matter in the “halo” of the galaxy to the core helium abundance at formation for stars in globular clusters, these astrophysical assumptions underlie the analyses of the effects of axions on observations and simulations. Turning these analyses on their head, we conclude that the physical picture, presented in this Letter, stimulates a new look at these conventional assumptions, which cut across a wide swath of astrophysical phenomena from star formation to supernovae, from galactic structure to galactic chemical evolution, and from the structure to the evolution of pulsars, white dwarfs, red giants, and horizontal branch stars (See Section 9).
2 0.5 eV QCD Axion Physical Picture
@@
The physical picture for dark matter, dark energy, and the Big Bang begins with an observation by Y. B. Zel’Dovich (1967). Stimulated by some of the first quasar absorption line spectroscopy (G. Burbidge, 1967), a number of astrophysicists suggested a value for the cosmological constant $\Lambda\approx+5\times 10^{-56}~{\rm cm}^{-2}$ (N. Kardashev, 1967; J. Shklovsky, 1967; V. Petrosian et al., 1967). In this context, Y. B. Zel’Dovich (1967) observed that the gravitational binding energy, $U_{\Lambda}$, generated by a massless particle and antiparticle, each with the kinetic energy, $E$, separated by the distance, $R=\hbar c/E$,
@@
$f_{A}$ $=$ $G\frac{\left(\hat{E}/c^{2}\right)^{2}}{\hbar c/\hat{E}}$ (13)
$=$ $\frac{\hat{E}^{3}}{E_{\rm Pl}^{2}}.$
@@
3 0.5 eV QCD Axion Cosmological Parameters
Within the physical picture sketched in Section 2 (See also Appendix A), the most straightforward cosmological parameter to estimate is that of the 0.5 eV QCD axion dark matter, $\Omega_{A}h^{2}$, with the energy density, $u_{A}=n_{A}m_{A}c^{2}$, given by the axion rest-mass energy, $m_{A}c^{2}$, and by the axion number density, $n_{A}$:
@@
where the axion-to-photon number ratio, $n_{A}/n_{\gamma}=6$, is set to its value at the Big Bang within the physical picture sketched in Section 2 (For the derivation, see Appendix A). Combining this value for $n_{A}$ with the value of $m_{A}c^{2}$ from Eq. (6), we find a high-precision prediction from first principles for the cosmological parameter of the cold dark matter in the universe,
@@
where the exact value of the astronomical unit ${\rm AU}=149~597~870.700~{\rm km}$ (N. Capitaine et al., 2012) was used to derive the numerically exact value of the cosmological time, $h/H_{0}=({\rm AU}/100~{\rm km})(180/\pi)(60)^{2}\times 10^{6}~{\rm sec}=3.085~677~581\times 10^{17}~{\rm sec}.$
As noted in Section 1, the high-precision prediction from first principles for the cold dark matter cosmological parameter in Eq. (18) agrees to better than a percent with the best present estimate for this parameter within the standard $\Lambda{\rm CDM}$ cosmology using early-universe observations (Planck Collaboration et al., 2019, 2020)
@@
Building on this apparent success of our physical picture for dark matter, dark energy, and the Big Bang in bridging the gap between early- and late-universe observations, a quick and dirty estimate of the cosmological constant from first principles is a natural next step. Combining the two estimates of the kinetic energy of radiation at the Big Bang, $\check{E}$, from Eq. (15) and Eq. (16), and using the high-precision first principles prediction for the value of the present expansion rate from Eq. (1), we get a percent-level precise estimate for the cosmological parameter, $\Omega_{\Lambda}$, for the cosmological constant
@@
Although necessarily rough, the estimate for the cosmological parameter of the cosmological constant in Eq. (20) still agrees to within twenty percent with the best present estimate for this parameter within the standard $\Lambda{\rm CDM}$ cosmology using early-universe observations (Planck Collaboration et al., 2019, 2020)
@@
To push the precision of the first principles estimate of the cosmological constant and to determine the last remaining cosmological parameter, $\Omega_{B}h^{2}$, that of the baryons, we must leverage the numerically exact value for the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}=0.399~286~(14)~[35~{\rm ppm}]$, given in Eq. (5), and derived in Sec. 4:
@@
where the value predicted within our physical picture in Eq. (18) was used for $\Omega_{A}h^{2}$. As expected, this high-precision first principles prediction for the baryon abundance in Eq. (22) differs sharply, by $190\sigma$, roughly a factor of two, from the best present estimate for this parameter within the standard $\Lambda{\rm CDM}$ cosmology using early-universe observations (Planck Collaboration et al., 2019, 2020)
@@
As discussed in Section 1 (See Eq. (4)), this sharp discrepancy in the value for the baryon abundance predicted with our physical picture and the value predicted with the standard $\Lambda{\rm CDM}$ cosmology is the key to resolving the Hubble tension between early- and late-universe observations, since the values of the other cosmological parameters are basically the same in our physical picture as in $\Lambda{\rm CDM}$ cosmology.
@@
where the value of $\sqrt{\chi_{\rm QCD}}$ agrees with the best present estimate from chiral perturbation theory (CPT) $\sqrt{\chi_{\rm QCD}}({\rm CPT})=5~690~(50)~{\rm MeV}^{2}~[8.8~{\rm per~mille}]$ (M. Gorghetto & G. Villadoro, 2019), to within two percent, but is nearly 26000 times more precise. At last, combining the Big Bang temperature from Eq. (3) with the 0.5 eV QCD axion rest-mass energy from Eq. (6) yields the cosmological parameter for the cosmological constant, $\Omega_{\Lambda}h^{2}$, to 22 ppm (See Appendix B for the derivation of the precision of the estimate and Appendix D for the derivation of the estimate itself):
@@
4 0.5 eV QCD Axion-to-Baryon Energy Density Ratio
To derive the numerically exact value of the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}$, given in Eq. (5), which is the key to the proposed resolution of the Hubble Tension, presented in Sections 1, 2, and 3, we use the cyclic cosmology matching condition (See Appendix C):
@@
where the axion energy density in the present universe, $u_{A}=m_{A}c^{2}n_{A}$, was used in the last line, and the value of the dimensionless constant, $N_{U}=1/(2\pi^{2})$, is derived in Appendix A. Combining the cyclic cosmology matching condition in Eq. (29) with the expression for the cosmological constant in terms of the dark energy density at the Big Bang in Eq. (30), we derive the remarkable, numerically exact baryon-to-axion ratio for energy density in the present universe given in Eq. (5)

### v49

@@ I Introduction
In light of the apparent success of 0.5 eV QCD axion dark matter in resolving the Hubble tension, these astrophysical “constraints” and “exclusions” raise important questions about the conventional assumptions for a wide range of astrophysical phenomena. From the velocity dispersion of dark matter in the “halo” of the galaxy to the core helium abundance at formation for stars in globular clusters, these astrophysical assumptions underlie the analyses of the effects of axions on observations and simulations. Turning these analyses on their head, we conclude that the physical picture, presented in this Letter, stimulates a new look at these conventional assumptions, which cut across a wide swath of astrophysical phenomena from star formation to supernovae, from galactic structure to galactic chemical evolution, and from the structure to the evolution of pulsars, white dwarfs, red giants, and horizontal branch stars (See Section IX).
II 0.5 eV QCD Axion Structure Formation
Figure 2: Cooling of 0.5 eV QCD axion dark matter in the early universe shows a cross-over in the adiabatic compressibility, $dp_{A}/du_{A}$ (violet pluses, +), from the radiation-like value, $dp_{A}/du_{A}\approx 1/3$, at large redshift, to the matter-like value, $dp_{A}/du_{A}\approx 0$, at small redshift, starting right around the redshift of “matter-radiation equality,” $z_{\rm eq}=3~376~(19)$ (vertical orange arrow), predicted by the standard cold dark matter cosmology with cosmological constant ($\Lambda{\rm CDM}$) using observations of the early universe (Planck Collaboration et al., 2019, 2020).
The formation of large-scale astrophysical structures in the early universe is a key test of any proposal for the nature of dark matter, dark energy, and the Big Bang, including 0.5 eV QCD axion cosmology. Historically, the now-standard cold dark matter cosmology first emerged in the context of concerns about the relative lack of large-scale anisotropy in the cosmic microwave background (CMB), compared to expectations based on the then-standard baryon-based Big Bang cosmology (P. J. E. Peebles, 2020, pg. 301–302). In fact, Peebles first proposed a hot dark matter cosmology (P. J. E. Peebles, 1982a), before turning to cold dark matter (P. J. E. Peebles, 1982b), with the shared goal, in both cases, of “washing out” large-scale CMB anisotropy and, in this way, resolving the tension between the then-standard cosmology and early-universe observations by D. J. Fixsen et al. (1983).
As the 0.5 eV QCD axion dark matter equation of state crosses over to matter-like at small redshifts from radiation-like at large redshifts, starting around $z_{\rm eq}$, as shown in Fig. 2, the dynamics of the expansion change to “matter-dominated” from “radiation-dominated.” Indeed, with six axions for each photon made in the cosmic background radiation (See Section VII and Appendix A), axions dominate the expansion dynamics on both sides of “matter-radiation equality.” The cross-over in the axion equation of state (Fig. 2), beginning around $z_{\rm eq}$, drives the cross-over in the character of the expansion which then drives the large-scale structure formation seen by early-universe observations, just as the change from “radiation-domination” to “matter-domination” drives such structure to form in the standard cold dark matter cosmology (P. J. E. Peebles, 1993, pgs. 624–625).
The rough agreement, shown in Fig. 2, between the best present observational value of the redshift where large-scale structure formation begins, $z_{\rm eq}$, and the predicted value for this redshift based on the cross-over in the equation of state for the 0.5 eV QCD axion dark matter, is a crucial test for the proposed cosmology. Together with the striking agreement between the predicted and observed values for the present energy density of cold dark matter in the universe (See Section VII), this agreement on large-scale structure formation provides compelling evidence that the proposed 0.5 eV QCD axion dark matter cosmology is compatible with early-universe observations. In this way, we deepen the proposed resolution of the Hubble tension, sketched in Section I, by ensuring that the agreement between the standard $\Lambda{\rm CDM}$ cosmology and early-universe observations is not spoiled by the switch to “cooling” dark matter, in the form of 0.5 eV QCD axion dark matter (Fig. 2).
III 0.5 eV QCD Axion Nucleosynthesis
Figure 3: Cooling of the universe in the first three minutes after the Big Bang for a mix with six 0.5 eV QCD axions for each photon (violet line, ) and for the standard mix of photons and light leptons (filled and unfilled boxes) both reach the freeze-out of the neutron-to-proton ratio (red horizontal line with dots) around the same time, but the standard mix (P. J. E. Peebles, 1966; G. Steigman, 2007) takes roughly twice as long to start forming helium nuclei (green horizontal line with boxes).
Indeed, doubling the baryon number density, $n_{B}(T_{d})$, compared to the standard cold dark matter cosmology value, as we propose to do to resolve the Hubble tension (See Section I and VII), ensures a nearly complete conversion of deuterons into helium, and, thus, a nearly complete absence of primordial deuterium (R. J. Cooke et al., 2018), for the timing of the turn-off, $t_{d}(\rm{\Lambda CDM})\approx 200~{\rm sec}$ (P. J. E. Peebles, 1966; G. Steigman, 2007), set by the standard mix of light leptons and photons. Yet, with the appropriate mix of six axions for every photon made in the Big Bang (See Section VII and Appendix A), the turn-off time, $t_{d}\approx 100~{\rm sec}$, comes nearly twice as fast as it does with the standard mix (See Fig. 3). Thus, the value of the product that determines the deuterium abundance, $n_{B}(T_{d})\langle\sigma v\rangle t_{d}\sim 1$ (P. J. E. Peebles, 1993, pg. 141–142), lands in about the same place, preserving the agreement with observations, where $\langle\sigma v\rangle$ is the thermal average of the product of the neutron capture cross-section, $\sigma$, with the nucleon relative velocity, $v$.
Still, the change in the time of the turn-off, $t_{d}$, has observable consequences due to the change in the gap in time, $t_{d}-t_{n}$, between the start of BBN, at $t_{d}$, and the “freeze-out,” at $t_{n}$, of the weak nuclear reactions between light leptons and nucleons (Fig. 3). During this gap, the decay of free neutrons with life-time, $\tau_{n}\approx 878~{\rm sec}$ (P. J. Mohr et al., 2025), cuts down the neutron-to-proton ratio (P. J. E. Peebles, 1993, pg. 184–186). By cutting $t_{d}$ in half, while leaving $t_{n}$ more or less unchanged, we shorten the gap, $t_{d}-t_{n}$, by about ten percent of the free neutron lifetime, $\tau_{n}$, so that about ten per cent more neutrons make it to the start of BBN, at $t_{d}$, after which essentially all the surviving free neutrons end up as primordial helium (P. J. E. Peebles, 2020, pg. 128–129): We thus expect a roughly ten percent larger value for the primoridal helium abundance, $Y_{\rm P}\approx 0.27$, compared to the best present estimate with the standard mix of light leptons and photons, $Y_{\rm P}({\rm\Lambda CDM})\approx 0.25$ (C. Pitrou et al., 2018; S. Gariazzo et al., 2022).
IV 0.5 eV QCD Axion Initial Fluctuations
More than five decades ago, E. R. Harrison (1970) and Y. B. Zeldovich (1972) argued for a scale-invariant initial spectrum of fluctuations in the energy density of the radiation created in the Big Bang, $P(k)=Ak^{n_{s}}$, where the spectral index value, $n_{s}=1$, ensures scale-invariance, and the constant, $A$, is independent of the co-moving wavevector of the fluctuation, $k$. Within the standard $\Lambda{\rm CDM}$ cosmology, the best present estimate from early-universe observations for the value of the spectral index, $n_{s}({\rm\Lambda CDM})=0.9677~(37)$ (Planck Collaboration et al., 2019, 2020), falls $8.8\sigma$ below the scale-invariant value, $n_{s}=1$. By contrast, the conformal symmetry of the cross-over from future infinity to the next big bang in 0.5 eV QCD axion cosmology (See Appendix C) requires the scale-invariant value, $n_{s}=1$, in agreement with the argument of E. R. Harrison (1970) and Y. B. Zeldovich (1972), but in sharp tension ($8.8\sigma$) with the standard $\Lambda{\rm CDM}$ value.
V 0.5 eV QCD Axion Astrophysical Constraints
Figure 4: Axion-photon coupling grows with initial core helium mass fraction in simulations of stellar evolution for horizontal branch stars constrained by observations of globular clusters (curved blue lines with dots) (A. Ayala et al., 2014). The predicted value for the axion-photon coupling strength (horizontal green lines with boxes), given in Eq. (10), combines with this empirical-computational relation to give a prediction for the helium mass fraction (red box with whiskers) that differs by $5.0\sigma$ from the primordial helium abundance expected within the standard cold dark matter cosmology with cosmological constant ($\Lambda{\rm CDM}$) based on simulations of Big Bang nucleosynthesis (thick vertical orange line) (S. Gariazzo et al., 2022; C. Pitrou et al., 2018), but that agrees rather well with the best present estimates for the initial core helium mass fraction in the Sun (pair of thin vertical red lines) (L. Piersanti et al., 2007; A. M. Serenelli & S. Basu, 2010).
$R=6.26Y_{\rm HB}-0.41g_{10}^{2}-0.12,$ (11)
where $R=N_{\rm HB}/N_{\rm RGB}=1.39~(3)$ was the observed ratio of the number of HB stars, $N_{\rm HB}$, to the number of stars on the red giant branch (RGB), $N_{\rm RGB}$, of the color-magnitude diagram in 39 globular clusters. Combining the empirical-computational relation in Eq. (11) with the predicted value for the axion-photon coupling in Eq. (10), we find a percent-level precise estimate for the initial core helium mass fraction in these HB stars that formed in the first billion years after the Big Bang:
$Y_{\rm HB}=0.272~(5)~[18~{\rm per~mille}],$ (12)
which differs by $5\sigma$ with the best present estimate of the “primordial” helium abundance, $Y_{P}({\rm\Lambda CDM})=0.247~09~(17)~[0.69~{\rm per~mille}]$ (C. Pitrou et al., 2018; S. Gariazzo et al., 2022), within the standard cold dark matter cosmology, but that roughly matches the best present estimates for the initial core helium mass fraction in the Sun, $Y_{\odot}=0.273~(6)~[{\rm 22~per~mille}]$ (A. M. Serenelli & S. Basu, 2010; L. Piersanti et al., 2007). As discussed in Section III, the mix of axions and photons that dominates the expansion of the universe leading up to Big Bang nucleosynthesis, within the proposed physical picture, yields about ten percent more primordial helium, $Y_{P}\approx 0.27$, than the standard mix of light leptons and photons, and, in this way, resolves the $5\sigma$ tension between $Y_{\rm HB}$ in Eq. (12) and $Y_{P}({\rm\Lambda CDM})$.
Similarly, the recent attempts to “rule out” axion dark matter with the rest-mass energy in Eq. (9) and the axion-photon coupling in Eq. (10) using “blank-sky” observations of the dark matter “halo” of the Milky Way at the near-infrared wavelength $2\lambda_{A}=2h/m_{A}c\approx 4.92~{\rm\mu m}$, corresponding to the decay of the axion into a pair of photons, subtract a “smooth astrophysical background” by fitting a quartic polynomial to the near-infrared spectrum within a window of width $\Delta\lambda\approx 0.2~{\rm\mu m}$ (E. Pinetti, 2025). However, the proposed physical picture implies that 0.5 eV QCD axion dark matter is in thermal equilibrium with the cosmic microwave background (See Section VII and Appendix A), and, thus, the axions in the dark matter “halo” of the Milky Way have velocity dispersion, $\sigma_{v}$, fixed by the CMB temperature, $T_{\gamma}$, and by the 0.5 eV QCD axion dark matter rest-mass energy, $m_{A}c^{2}$, leading to an expected line-width, $2\Delta\lambda_{A}$, that is roughly the same size as the window used to subtract the “smooth astrophysical background”:
$2\Delta\lambda_{A}$ $=$ $2\lambda_{A}\frac{\sigma_{v}}{c}=2\lambda_{A}\sqrt{\frac{3kT_{\gamma}}{m_{A}c^{2}}}\approx 0.2~{\rm\mu m}.$ (13)
VI 0.5 eV QCD Axion Physical Picture
@@
The physical picture for dark matter, dark energy, and the Big Bang begins with an observation by Y. B. Zel’Dovich (1967). Stimulated by some of the first quasar absorption line spectroscopy (G. Burbidge, 1967), a number of astrophysicists suggested a value for the cosmological constant, $\Lambda\approx+5\times 10^{-56}~{\rm cm}^{-2}$ (N. Kardashev, 1967; J. Shklovsky, 1967; V. Petrosian et al., 1967). In this context, Y. B. Zel’Dovich (1967) observed that the gravitational binding energy, $U_{\Lambda}$, generated by a massless particle and antiparticle, each with the kinetic energy, $E$, separated by the distance, $R=\hbar c/E$,
@@
$f_{A}$ $=$ $G\frac{\left(\hat{E}/c^{2}\right)^{2}}{\hbar c/\hat{E}}=\frac{\hat{E}^{3}}{E_{\rm Pl}^{2}}.$ (19)
@@
VII 0.5 eV QCD Axion Cosmological Parameters
Within the physical picture sketched in Section VI (See also Appendix A), the most straightforward cosmological parameter to estimate is that of the 0.5 eV QCD axion dark matter, $\Omega_{A}h^{2}$, with the energy density, $u_{A}=n_{A}m_{A}c^{2}$, given by the axion rest-mass energy, $m_{A}c^{2}$, and by the axion number density, $n_{A}$:
@@
where the axion-to-photon number ratio, $n_{A}/n_{\gamma}=6$, is set to its value at the Big Bang within the physical picture sketched in Section VI (For the derivation, see Appendix A). Combining this value for $n_{A}$ with the value of $m_{A}c^{2}$ from Eq. (9), we find a high-precision prediction from first principles for the cosmological parameter of the cold dark matter in the universe,
@@
where the exact value of the astronomical unit, ${\rm AU}=149~597~870.700~{\rm km}$ (N. Capitaine et al., 2012), was used to derive the numerically exact value of the cosmological time, $h/H_{0}=({\rm AU}/100~{\rm km})(180/\pi)(60)^{2}\times 10^{6}~{\rm sec}=3.085~677~581\times 10^{17}~{\rm sec}.$
As noted in Section I, the high-precision prediction from first principles for the cold dark matter cosmological parameter in Eq. (24) agrees to better than a percent with the best present estimate for this parameter within the standard $\Lambda{\rm CDM}$ cosmology using early-universe observations (Planck Collaboration et al., 2019, 2020):
@@
Building on this apparent success of our physical picture for dark matter, dark energy, and the Big Bang in bridging the gap between early- and late-universe observations, a quick and dirty estimate of the cosmological constant from first principles is a natural next step. Combining the two estimates of the kinetic energy of radiation at the Big Bang, $\check{E}$, from Eq. (21) and Eq. (22), and using the high-precision first principles prediction for the value of the present expansion rate from Eq. (1), we get a percent-level precise estimate for the cosmological parameter, $\Omega_{\Lambda}$, for the cosmological constant,
@@
Although necessarily rough, the estimate for the cosmological parameter of the cosmological constant in Eq. (26) still agrees to within twenty percent with the best present estimate for this parameter within the standard $\Lambda{\rm CDM}$ cosmology using early-universe observations (Planck Collaboration et al., 2019, 2020):
@@
To push the precision of the first principles estimate of the cosmological constant and to determine the last remaining cosmological parameter, $\Omega_{B}h^{2}$, that of the baryons, we must leverage the numerically exact value for the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}=0.399~286~(14)~[35~{\rm ppm}]$, given in Eq. (7), and derived in Sec. VIII:
@@
where the value predicted within our physical picture in Eq. (24) was used for $\Omega_{A}h^{2}$. As expected, this high-precision first principles prediction for the baryon abundance in Eq. (28) differs sharply, by $190\sigma$, roughly a factor of two, from the best present estimate for this parameter within the standard $\Lambda{\rm CDM}$ cosmology using early-universe observations (Planck Collaboration et al., 2019, 2020):
@@
As discussed in Section I (See Eq. (4)), this sharp discrepancy in the value for the baryon abundance predicted with our physical picture and the value predicted with the standard $\Lambda{\rm CDM}$ cosmology is the key to resolving the Hubble tension between early- and late-universe observations, since the values of the other cosmological parameters are basically the same in our physical picture as in $\Lambda{\rm CDM}$ cosmology.
@@
where the value of $\sqrt{\chi_{\rm QCD}}$ agrees with the best present estimate from chiral perturbation theory (CPT) $\sqrt{\chi_{\rm QCD}}({\rm CPT})=5~690~(50)~{\rm MeV}^{2}~[8.8~{\rm per~mille}]$ (M. Gorghetto & G. Villadoro, 2019), to within two percent, but is nearly 26000 times more precise. At last, combining the Big Bang temperature from Eq. (VII) with the 0.5 eV QCD axion rest-mass energy from Eq. (9) yields the cosmological parameter for the cosmological constant, $\Omega_{\Lambda}h^{2}$, to 22 ppm (See Appendix B for the derivation of the precision of the estimate and Appendix D for the derivation of the estimate itself):
@@
VIII 0.5 eV QCD Axion-to-Baryon Energy Density Ratio
To derive the numerically exact value of the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}$, given in Eq. (7), which is the key to the proposed resolution of the Hubble Tension, presented in Sections I, VI, and VII, we use the cosmology matching condition (See Appendix C):
@@
where the axion energy density in the present universe, $u_{A}=m_{A}c^{2}n_{A}$, was used in the last line, and the value of the dimensionless constant, $N_{U}=1/(2\pi^{2})$, is derived in Appendix A. Combining the cosmology matching condition in Eq. (35) with the expression for the cosmological constant in terms of the dark energy density at the Big Bang in Eq. (36), we derive the remarkable, numerically exact baryon-to-axion ratio for energy density in the present universe given in Eq. (7)